Find the output needed to maximize profit given that the total cost and total revenue functions are TC=2Q and TR=100 ln (Q+1) respectively. 


1
Expert's answer
2022-03-25T10:29:54-0400

P=TR-TC=100ln(Q+1)-2Q

P'(Q)=100/(Q+1)-2=0

Q=100/2-1=49